OpenStudy (smurfy14):

[(x)/(x+2)]-[(2)/(2x-1)]=(1/5) how do you solve to get x=-3/4 as your answer??

OpenStudy (immanuelv):

smurfy14 i used this equation (2 x^2-3 x-4)/((x+2) (2 x-1)) = 1/5

OpenStudy (gw2011):

Multiply both sides by (x+2)(2x-1) to eliminate the denominator and you get: x(2x-1)-2(x+2)=1/5(x+2)(2x-1) Multiply both sides by 5 to eliminate 5 and you get: 5x(2x-1)-10(x+2)=(x+2)(2x-1) 10x^2-5x-10x-20=2x^2+3x-2 10x^2-15x-20=2x^2+3x-2 Now combine the like terms: 8x^2-18x-18=0 2(4x^2-9x-9)=0 Divide both sides by 2 to eliminate the 2 and you get: 4x^2-9x-9=0 (4x+3)(x-3)=0 4x+3=0 4x=-3 x=-3/4 x-3=0 x=3 Therefore, x=-3/4 and x=3

OpenStudy (smurfy14):

you're the best thanks so much!!

OpenStudy (gw2011):

You're welcome